mary earned $48000 from royalties on her book.she paid a 28% income tax on the royalties.she invested $15000 at one rate and the balance at a rate that was 1% lower,earning $1878 annual interest on the two investments. what was the higher interest rate?
Answer by Boreal(15235)

$48000*0.28=$13440 in tax, so $34,560 remaining
15000*x=15000x, where x=rate in decimal form
19560*(x-.01)=19560-195.6x
Those two equal 1878
Therefore, 15000x+19560x-195.6=1878
34560x-195.6=1878
34560x=2073.6
x=0.06 or 6% * 15000=$900
x-.01=0.05 or 5% *19560=$978
They add to $1878.